# Bridging the Gulf of Envisioning: Cognitive Design Challenges in LLM Interfaces

### v1

@@ 3.3. Recruiting New Pathways for Interacting With LLMs
In the Intentionality Gap, users interact directly with the LLM, predominantly in evaluation mode. This means they’re often trying to make judgments about the generated output and how changes to the input may produce different outputs aligned with their goal. With LLMs, they must attempt these cognitive maneuvers without a foundational context or a cohesive mental model of what the LLM is doing functionally. In a third proposed interaction pathway (Figure 2-3), we suggest a dedicated interface tool such as a Word Processor that can support the user in developing intentions and actions more directly tied to their task goals. This shift to an interface tool that knows about tasks – how to support writing and editing – facilitates the user’s induction of a function-based model of the task process they envision, enabling the user to more precisely calibrate and synchronize their intentions with corresponding action descriptions. Another alternative to this pathway is integrating LLMs within existing tools, as we will see in section 4. The use of a specialized interface tool can help scaffold humans interacting with LLMs in this context, providing structured prompts or templates to guide users in clarifying their intentions and fostering a clearer sense of content direction. This added structure acts as a scaffold to help users frame their output evaluations with a more informed perspective on what is required to generate a successful output, reducing the intentionality gap. Further, a specialized tool can guide users in specifying procedural aspects of their intentions, thereby reducing the need for repeated trial-and-error.
@@
As discussed earlier, a cognitive task model such as Hayes and Flowers (Hayes, 2013) describes how a writer moves from a goal to the executed text output. The steps of creating a story context and characters, and then planning, executing, and revising, describe the process of generation across writers. The steps for how to write a story are the subject of much analysis and instruction to help people become better writers. While this process is not the only viable way to generate a story, it is a way people often use, and therefore may be evident in the stories captured in the textual data corpus employed in training LLMs. While the LLM was not trained in this task process, it emerges through its analysis across words, phrases, and stories, and becomes implicitly embedded in the LLM’s generation performance. That is why it is helpful for a user to mentally envision what is required for them to perform the generative task: Doing so likely engages the same task knowledge used by other writers with similar intentions. Then, describing their intentions to the LLM with those cognitively-informed specifications makes accessing stories with similar intent more likely through the lexical indexes built through transformer networks.
@@
To traverse the extreme database of connections encoded within large learning models and generate a desired outcome, a user must attempt to describe not only what others wrote about, but also, why they wrote. Without understanding intentions, it is impossible to assess the actions leading to good outcomes in the past, and impossible to assess outcomes from LLM systems as meeting goals. The value of generative outputs depends entirely on the intentions linking goals to their execution. The three pathways approach shown in Figure 2 builds on human knowledge of successful task completion. The three gaps, including the capability gap, language gap, and intentionality gap, together comprise the gulf of envisioning.
@@ 4. Examining the Envisioning Gap in Three LLM Interfaces
Up until now, we have characterized the envisioning gulf based on the core capabilities of LLMs (i.e., language understanding and text generation) with a focus on writing tasks. We now analyze the design of three existing LLM interfaces (see Figure  3) across different generative tasks to pinpoint how the three gaps manifest during interactions. Concretely, we look at how LLM interface designs support the specific cognitive task processes they’re built for – using the framework of planning, execution, and evaluation – and the specific features they implement to minimize the three gaps in our revised interaction model.
Figure 3. Example affordances for how (A.) ChatGPT, (B.) Spellburst, and (C.) Cursor bridge the capability (top), language (middle), and intentionality (bottom) gaps.
@@ 4.1. Writing using ChatGPT
ChatGPT (Figure  3-A) is an LLM with a corresponding web-based interface developed by OpenAI (OpenAI, 2023a). Through the interactive chat interface, users can supply prompts and engage in dynamic conversations with the model. In the context of a writing task, ChatGPT can quickly produce drafts or outlines based on a given topic, elaborate on an outline, perform grammar and style checks, and paraphrase and synthesize. Naturally, the full spectrum of writing support it offers is diverse and dynamically evolving.
@@ 4.1.1. Capability Gap
To support users in understanding the action space, the ChatGPT landing page currently provides example prompts and tasks the model can perform. These affordances begin to reduce the capability gap during interactions. Yet, the tasks depicted often lack the granularity required for users to devise concrete plans. Further, the tool also allows users to start different “chats” for different lines of planning and execution cycles. However, the key cognitive issue in planning is determining how to break down their goals into specific, actionable steps with the LLM. Simply having different chats may help with the organization, but it does not close the gap in helping users know how to best to formulate their intentions.
4.1.2. Language Gap
Users mainly discover how LLMs interpret prompts within ChatGPT using trial and error. These include features such as regeneration, editing the original prompt, and managing different “chats” with ChatGPT on the left sidebar. However, ChatGPT does not provide a history of previous outputs, making it difficult to compare the quality of regeneration from run to run. If users cannot see how slight language changes affect outputs, they might struggle to learn from their linguistic adjustments to the prompt. In other words, the lack of feedback inhibits their ability to understand how language nuances influence the model’s interpretations, making the language gap even more pronounced. Separately, ChatGPT allows users to set “custom instructions,” which specify details and guidelines when engaging in a dialogue with the model. However, this requires users to foresee these intentions upfront before interacting with the model.
@@ 4.1.3. Intentionality Gap
The intentionality gap deals with the challenge users face when assessing the LLM-generated text because they bypassed the planning and execution processes. Custom instructions are the primary mode for aligning user values with model output, as they act somewhat as “base prompts.” There are two parts to defining custom instructions: what you want ChatGPT to know about you as well as how you would like ChatGPT to respond. The sum of these two halves mimics domain-specific prompts seen in other LLM-enabled systems, as the former asks the model to play a role in narrowing its scope while the latter helps steer the model’s generation. Overall, this feature ensures that users no longer need to specify such context when prompting, allowing them to focus on crafting a good prompt based on their specific goals and intentions. However, even with this alignment, users may still encounter challenges in assessing the output since they may lack a comprehensive mental model of the content. Finally, at evaluation time, users can ask the LLM questions about its previous inputs, regenerate a response from the LLM, or even directly edit an old prompt to see how a model changes its answer. These features aim to give users more clarity on the LLM’s outputs, offering ways to refine and adjust the content to better match their intentions.
@@ 4.2. Creative Coding using Spellburst
Next, we look at Spellburst (Figure  3-B), a creativity support tool developed as a research artifact (Angert et al., 2023). We selected this tool as it aims to support the artists’ “exploratory creative coding” workflow to address cognitive challenges in creative work. Specifically, they focus on bridging the artist’s creative intents expressed in natural language with the implementation of those intents as code using LLM.
@@ 4.2.1. Capability Gap
4.2.2. Language Gap
Spellburst also uses a variety of techniques to bridge the language gap. After the interface is prompted to generate a sketch, the model interleaves comments within the code of the sketch that point to the specific change in the code based on the user’s prompt. These comments enable users to see how the model’s inputs prompt its outputs. As an example: if a user modified a sketch to turn the sketch’s outline from black to blue, the model would output the sketch with a code comment in the location where the outline color was changed. There are also numerous ways to experiment and see how the model may interpret different prompts, including the duplication and modification of a sketch node or the creation of a new branch within the interface entirely, representing a new line of semantic and syntactic jumps. Finally, two semantic operators help make the model’s output more explainable by asking the LLM to describe its own output – diffing and extracting. Diffing allows users to see both syntactic (code-level) and semantic (prompt-based) differences between two sketches, while extracting enables users to ask the model questions about a sketch.
@@ 4.2.3. Intentionality Gap
Based on their provided prompts, Spellburst asks the LLM to “act as an expert creative coder.” Through this, they are setting a high-level intentionality of how the LLM should be thinking about generating the output. Furthermore, the prompts help ensure that the model generates code that runs and compiles correctly. The team behind the system also crowd-sourced a set of image transformations to develop a creative coding taxonomy. This taxonomy drives the autocomplete suggestions, as users can orient their intentions alongside this shared vocabulary when prompting the model, improving the quality of the LLM’s output.
@@ 4.3. Software Development using Cursor
Cursor (Figure  3-C) is a code editor that helps users pair program with AI (Anysphere, 2023). The system uses an LLM – users can choose between GPT-4 (OpenAI, 2023b) or ChatGPT (OpenAI, 2023a) – to help developers chat with their project, make code changes, and address bugs. Given the growing interest in utilizing AI to supercharge the software development process (Github, 2023; Imai, 2022), we wanted to see how Cursor implemented a human-LLM interface for the complex cognitive task process of programming.
As a fork of VSCode (Microsoft, 2023), Cursor has numerous tool affordances critical to supporting programming, including a file explorer, line completion, and a search bar. However, there are many features specific to Cursor. When planning, users are given several examples of tasks they can accomplish. In addition, there are various places to converse with the LLM, including when selecting the text, using keyboard shortcuts within the main editor and the right sidebar dedicated to chatting with the model. During execution, the user can either write code normally in the editor or prompt the model to generate code. Finally, the system provides several features for evaluation, including regenerating model responses, chatting with a model about code, and auto-debugging.
@@ 4.3.1. Capability Gap
Cursor’s first way of showcasing what tasks are available within the system is through a set of example files when you first open the editor. Each of these files is named after a possible task, and the comments of the file give instructions regarding what keyboard shortcuts and prompt to use to complete the goal. Similar to ChatGPT, though, since these tasks are high-level (i.e., “fix a bug,” “explain code”), they are not helpful for more specific planning of goals and intentions. When interacting with code across the various panels of the interface – including the code editor and the chat interface on the right-hand side – the tooltip shows both what actions can be taken with an LLM as well as how to establish context. For instance, highlighting text within a code editor gives users the option to either add the code to the right hand chat interface or edit directly with a prompt. Likewise, inside of the chat interface, buttons appear for users to reference a file or change the scope of the user’s question within a prompt.
4.3.2. Language Gap
To help users understand how LLMs interpret language, Cursor allows users to ask models questions about their codebase. Topics for these questions can range from the inner workings of a particular function to the overall flow of an entire project. Since this feature can give answers about model-generated and user-written code, developers can get a better idea of how LLMs interpret their programs under the hood. A similar feature that can help decipher how models understand code is by prompting the model to generate inline comments, which is not done by default. Finally, despite there being no direct functionality to regenerate output outside of feeding the model the same prompt, there is an option to “rerun without context” to see the effects of adding code or references to other materials alongside a prompt. The lack of regeneration history, however, makes it more complicated for users to map changes in their inputs to changes in model outputs.
@@ 4.3.3. Intentionality Gap
The primary way Cursor aligns user intent with model output is through references of code snippets, files, and documentation. When authoring a prompt, users can attach an existing file of code they’ve written, a function within the code editor, or even documentation of third-party libraries to steer the output of the model. This functionality has many use cases, including prompting the model to adopt the same style or asking the LLM to use a particular method or function from a library. While this feature reduces the amount of effort involved in establishing context with the LLM, this affordance does not inherently give users a mental model of how the LLM works, meaning users may still run into issues. Furthermore, to help evaluate output, user can ask Cursor questions about their codebase. Since these questions can be about any piece of code within the code editor – either model- or user-generated – this feature can help users understand how to make their prompts more optimal.
@@ 5. Recommendations for Designing Interfaces for Envisioning
In addition to the three interfaces described in detail in the previous section, we conducted a qualitative analysis of 12 systems to identify design patterns that would potentially support the process of envisioning. We only selected tools that were either directly accessible or at least had a video demonstration and accompanying technical description of the features, such as from research papers. For each tool, we identified affordances along the three main operators in generative tasks, namely planning, execution, and evaluation. We then clustered the identified features based on their functional similarity to develop a set of interface design patterns and corresponding tenets for teams building human-LLM systems. We do not claim that our categorization is comprehensive but it provides a starting point for the design of interface affordances for envisioning LLM interactions.
@@
Design Pattern 1 – Visually Track Prompts and Outputs: LLM interactions are often iterative and require a trial-and-error approach. We observed that some tools provided users with a visual interface for capturing their prompting and divergent thinking pathways. A popular choice is node-based interfaces, which allow users to visualize multiple different outputs (nodes) and trace how they are connected (edges). As mentioned before, in Spellburst (Angert et al., 2023), each node represents a sketch, and edges showcase different iterations of the sketch through merging, diffing, and other semantic operators. In PromptChainer (Wu et al., 2022a), a system that helps users chain together LLM prompts for complex tasks, each node represents an individual prompt and corresponding output while each edge represents the use of the output as context for another prompt. Such external representations of their thought process not only lower the cognitive load, but users are better able to engage in more deliberate prompt authoring. For instance, users can readily see if their previous line of thinking was fruitful or make adjustments to subsequent prompts based on prior outputs. These affordances correspond to the tenet that: Users need guidance navigating the cognitive task space for prompt authoring, discerning which paths lead to desired or poor outcomes [T1].
@@
Design Pattern 2 – Suggest Ideas for Prompting: Many systems proactively offer users prompt suggestions. Some of them are aimed at assisting users who may not be as familiar with LLM-enabled interfaces, others serve as ideation partners in the cognitive task workflow. In addition to providing ideas for prompts, the examples and suggestions emphasize the importance of clarity and precision in language. If these suggestions come from examples the model is fine-tuned on, they can also help better align a user’s intentions with model behavior. For example, ChatGPT (OpenAI, 2023a) provides standalone prompt examples to showcase its utility. On the other hand, Notion (Notion, 2023), which is a tool for knowledge management, not only offers possible ways to use AI to improve your writing (i.e., “Change tone”): it also gives suggestions on how to prompt an LLM to steer these improvements in a certain direction (i.e., “Professional,” “Casual,” “Straightforward”). These patterns lead to two Tenets, namely: LLMs can serve as cognitive partners in task formulation [T2] , and A focus on clear, precise written language will help to bridge the gap between human intention and LLM output [T3].
@@
Design Pattern 3 - Provide Multiple Outputs: Rather than generate one example based on a user prompt, systems may provide numerous outputs from an LLM. This can be achieved by setting the temperature of the model – a parameter that dictates randomness – greater than $0$ and giving the model the same prompt or by explicitly asking the model to give more than one example. This feature allows users to view multiple options to see which one best fits their intentions. Furthermore, providing numerous outputs helps users link the effects of changes in prompts to changes in the final output of the model. Some systems also support grouping and clustering model outputs to make this process easier. For instance, BotDesigner (Zamfirescu-Pereira et al., 2023) lets users manually assign a tag to model outputs, while Sensecape (Suh et al., 2023) groups relevant topics together semantically based on a high-level topic. These features support the tenet that LLMs should communicate to users their divergent thinking strategies [T4].
@@
Design Pattern 4 - Make the Output Explainable: Some systems prompt LLMs to explain their outputs or to make them more interpretable. This design pattern allows users to better understand how LLMs interpret certain prompts and makes model outputs easier to use for manual editing. How this technique is applied in practice can differ depending on the domain task. Replit (Replit, 2023), a browser-based code editor, has an AI assistant named Ghostwriter that generates in-line comments within its code responses. Another code editor, Cursor (Anysphere, 2023), does not always provide code comments but does allow users to ask LLMs about the code they generate. In contrast, Sensecape (Suh et al., 2023), which is designed for exploration and sensemaking, prompts an LLM to return a response at different levels of detail, such as through summaries and keywords. These features help users address their intentionality gaps and better assess the model output. This pattern supports the tenet An error in human-LLM interaction is not only a user’s mistake or LLM failure but a breakdown in the distributed cognitive system. Envisioning repair should be collaboratively accomplished [T5].
@@
Design Pattern 5 - Use domain-specific prompting strategies: Outside of standard prompt engineering techniques, most systems use a custom prompting strategy depending on their task. These methods help steer the outputs from LLMs into something usable for the end goal while also minimizing the output ambiguity that may come from trying different prompts. As an example: Graphologue (Jiang et al., 2023), which is designed to turn text-based responses from LLMs into graphical diagrams, uses prompting techniques to have the models annotate entities and relationships within their outputs to create the diagrams in real time. Coding Steps (Kazemitabaar et al., 2023), a web-based application to help novices learn Python, prompts models with static examples, then user code, then the user prompt, to ensure that the level of output is appropriate for beginners. These strategies allow designers to implement the task conceptual model for the users and consequently allow users to assume task-specific system mental models. The corresponding tenet is that Users would favor having a system mental model and actions [T6].
@@
Design Pattern 6 - Allow for manual control of the output Many systems afford users the opportunity to manually edit the outputs and interactions with LLMs. Since a lot of LLM-enabled systems are built for exploration and ideation, direct manipulation can help users better incorporate their values and intentions into the model. Oftentimes, manual editing is introduced when one output is used as input to another LLM. For instance, while LIDA (Dibia, 2023), a tool for generating visualizations and infographics, prompts an LLM to output goals for dataset exploration, users are also allowed to enter their own goals and adjust the model’s suggestions. Likewise, Mirror (Xu et al., 2023) – an NLI for data querying and summarization – gives users the ability to edit the SQL queries generated by a pre-trained code model to add human expertise. These features align with the tenet If tasks are well-defined, people prefer dedicated interfaces over dynamic interfaces [T7].

### v2

@@ 3.3. Recruiting New Pathways for Interacting With LLMs
In the Intentionality Gap, users interact directly with the LLM in evaluation mode. This means they’re often trying to make judgments about the generated output and whether changes to the input may produce different outputs more aligned with their goal. With LLMs, they must attempt these cognitive maneuvers without a foundational context or a cohesive mental model of the LLM functions. In a third proposed interaction pathway (Figure 3- PATH 3), we suggest a dedicated interface tool with support for the user to further develop intentions and actions to more clearly direct the LLM toward their task goals. This shift to an interface tool that knows specific complex generative tasks – how to write a screenplay or code in Java – facilitates the user’s identification of a functional basis for the task process, enabling the user to more precisely specify, calibrate, and synchronize their intentions with corresponding plan descriptions. Critically, the scaffolding provided by such a “prompt prompter” can facilitate prompt entry and completeness by identifying expected elements of task functions that are missing in a user prompt and requesting needed information (e.g., “who is your audience?”). Another alternative pathway is integrating LLMs within existing functional tools, as we will see in section 4. The use of a specialized interface tool can help scaffold human interactions with LLMs by dynamically providing structured templates to guide users in clarifying their intentions and fostering clearer directions. This added structure serves as a scaffold for helping users frame their output evaluations with a more informed perspective on what is required in a prompt to generate a successful output, reducing the intentionality gap. While expert users can develop their own cognitive task models through LLM use to identify their own scaffolds, sharing that expertise can greatly help occasional and novice users. Furthermore, a specialized tool can guide users in specifying procedural aspects of their intentions, thereby reducing the need for repeated trial-and-error. Importantly, this prompting scaffold can change dynamically within an LLM session to better reflect the intentions as the user develops them.
@@
As discussed earlier, a cognitive task model such as Hayes and Flowers (Hayes, 2013) describes how a writer moves from a goal to the executed text output. While this process is not the only viable way to generate a story, it is an approach that people often use and, therefore, may be evident in the stories captured in the textual data corpus employed in training LLMs. While the LLM was not trained to write stories, this underlying cognitive model becomes implicitly embedded in the LLM through its training set of words, phrases, and stories, and emerges in the LLM’s generation performance. While the LLM does not have an explicit cognitive model of how to generate a story, it is built through statistical analysis of a story corpus created by human writers. That is why it is helpful for a user to mentally envision what is required for them to perform the generative task: Doing so likely engages the same task knowledge used by other writers with similar intentions. Then, describing their developed intentions to the LLM using those cognitively informed specifications makes accessing stories with similar intent more likely through the lexical indexes built into transformer networks.
@@
To traverse the extreme database of connections encoded within large learning models and generate a desired outcome, a user must attempt to describe not only what others wrote about, but also, why they wrote. Without understanding intentions well, it is impossible to access plans from past tasks leading to good outcomes, and impossible to assess outcomes from LLM systems as meeting one’s goals. The value of generative outputs depends entirely on the intentions linking goals to their execution in output. The three pathways shown in Figure 3 build on human knowledge of successful task completion and vary in the effort and process required. The three gaps, including the capability gap, instruction gap, and intentionality gap, together comprise the gulf of envisioning.
@@ 4. Examining the Envisioning Gap in Three LLM Interfaces
Up until now, we have characterized the envisioning gulf based on the core capabilities of LLMs (i.e., language understanding and text generation) with a focus on writing tasks. We now analyze the design of three existing LLM interfaces (see Figure  4) across different generative tasks to pinpoint how the three gaps manifest during interactions. Concretely, we look at how LLM interface designs support the specific cognitive task processes they’re built for – using the framework of planning, execution, and evaluation – and the specific features they implement to minimize the three gaps in our revised interaction model.
@@ 4.1. Writing using ChatGPT
ChatGPT (Figure  4-A) is an LLM with a corresponding web-based interface developed by OpenAI (OpenAI, 2023a). Through the interactive chat interface, users can supply prompts and engage in dynamic conversations with the model. In the context of a writing task, ChatGPT can quickly produce drafts or outlines based on a given topic, elaborate on an outline, perform grammar and style checks, and paraphrase and synthesize. Naturally, the full spectrum of writing support it offers is diverse and dynamically evolving.
@@ 4.1.1. Capability Gap
To support users in understanding the action space, the ChatGPT landing page currently provides example prompts and tasks the model can perform. These affordances begin to reduce the capability gap during interactions. Yet the tasks depicted often lack the granularity required for users to devise concrete plans. Further, the tool also allows users to start different “chats” for different lines of planning and execution cycles. However, the key cognitive issue in planning is determining how to break down their goals into specific, actionable steps with the LLM. Simply having different chats may help with the organization, but it does not close the gap in helping users know how to best to formulate their intentions.
4.1.2. Instruction gap
Users mainly discover how LLMs interpret prompts within ChatGPT through trial and error. These trial and error actions subsume features such as regeneration, editing the original prompt, and managing different “chats” with ChatGPT on the left sidebar. Yet ChatGPT does not provide a history of previous outputs, making it difficult to compare the quality of regeneration from run to run. If users cannot see how slight language changes affect outputs, they might struggle to learn from their linguistic adjustments to the prompt. In other words, the lack of feedback inhibits their ability to understand how language nuances influence the model’s interpretations, making the Instruction gap even more pronounced. Separately, ChatGPT allows users to set “custom instructions,” which specify details and guidelines when engaging in a dialogue with the model. However, this feature requires users to foresee these intentions upfront before interacting with the model and a more general understanding of instruction utility across tasks.
@@ 4.1.3. Intentionality Gap
The intentionality gap reflects the challenge users face when evaluating the LLM-generated text because they bypassed the planning and execution processes. Custom instructions are the primary mode for aligning user values with model output, acting somewhat as “base prompts.” There are two parts to defining custom instructions: (1) what you want ChatGPT to know about you as well as (2) how you would like ChatGPT to respond. The sum of these two mimics domain-specific prompts seen in other LLM-enabled systems, as the former asks the model to play a role in narrowing its scope while the latter helps steer the model’s generation. Overall, this feature ensures that users no longer need to specify such context when prompting, allowing them to focus on crafting a good prompt based on their specific goals and intentions. However, even with this alignment, users may still encounter challenges in assessing the output because they may lack a comprehensive mental model of the content. Finally, at evaluation time, users can ask the LLM questions about its previous inputs, regenerate a response from the LLM, or even directly edit an old prompt to see how a model changes its answer. These features aim to give users more clarity on the LLM’s outputs, offering ways to refine and adjust the content to better match their intentions.
Figure 4. Example affordances for how (A.) ChatGPT (OpenAI, 2023a), (B.) Spellburst (Angert et al., 2023), and (C.) Cursor (Anysphere, 2023) bridges the capability (top), language (middle), and intentionality (bottom) gaps.
@@ 4.2. Creative Coding using Spellburst
Next, we look at Spellburst (Figure  4-B), a creativity support tool developed as a research artifact (Angert et al., 2023). We selected this tool as it aims to support the artists’ “exploratory creative coding” workflow to address cognitive challenges in creative work. Specifically, they focus on bridging the artist’s creative intents expressed in natural language with the implementation of those intents as code using LLM.
@@ 4.2.1. Capability Gap
4.2.2. Instruction gap
Spellburst also uses a variety of techniques to bridge the Instruction gap. After the interface is prompted to generate a sketch, the model interleaves comments within the code of the sketch that point to the specific change in the code based on the user’s prompt. These comments enable users to see how the model’s inputs prompt its outputs. As an example, if a user modified a sketch to turn its outline from black to blue, the model would output the sketch with a code comment in the location where the outline color was changed. There are also numerous ways to experiment and see how the model may interpret different prompts, including the duplication and modification of a sketch node or the creation of a new branch within the interface entirely, representing a new line of semantic and syntactic jumps. Finally, two semantic operators help make the model’s output more explainable by asking the LLM to describe its own output – diffing and extracting. Diffing allows users to see both syntactic (code-level) and semantic (prompt-based) differences between two sketches, while extracting enables users to ask the model questions about a sketch.
@@ 4.2.3. Intentionality Gap
Based on their provided prompts, Spellburst asks the LLM to “act as an expert creative coder.” Through this approach, they are setting a high-level intentionality of how the LLM should be thinking about generating the output. Furthermore, the prompts help ensure that the model generates code that runs and compiles correctly. The team behind the system also crowd-sourced a set of image transformations to develop a creative coding taxonomy. This taxonomy drives the autocomplete suggestions, as users can orient their intentions alongside this shared vocabulary when prompting the model, improving the quality of the LLM’s output.
@@ 4.3. Software Development using Cursor
Cursor (Figure  4-C) is a code editor that helps users pair program with AI (Anysphere, 2023). The system uses an LLM – users can choose between GPT-4 (OpenAI, 2023b) or ChatGPT (OpenAI, 2023a) – to help developers chat with their project, make code changes, and address bugs. Given the growing interest in utilizing AI to supercharge the software development process (Github, 2023; Imai, 2022), we analyze how Cursor implemented a human-LLM interface for the complex cognitive task process of programming.
As a fork of VSCode (Microsoft, 2023), Cursor has numerous tool affordances critical to supporting programming, including a file explorer, line completion, and a search bar. However, there are many features specific to Cursor. When planning, users are given several examples of tasks they can accomplish. In addition, there are various ways to converse with the LLM, including when selecting the text, using keyboard shortcuts within the main editor, and the right sidebar dedicated to chatting with the model. During execution, the user can either write code normally in the editor or prompt the model to generate code. Finally, the system provides several features for evaluation, including regenerating model responses, chatting with a model about code, and auto-debugging.
@@ 4.3.1. Capability Gap
Cursor’s first way of showcasing what tasks are available within the system is through a set of example files when a user first opens the editor. Each of these files is named after a possible task, and the comments in the file give instructions regarding what keyboard shortcuts and prompt to use to complete the goal. Similar to ChatGPT, though, since these tasks are high-level (i.e., “fix a bug,” “explain code”), they are not helpful for more specific planning of goals and intentions. When interacting with code across the various panels of the interface – including the code editor and the chat interface on the right-hand side – the tooltip shows both what actions can be taken with an LLM and how to establish context. For instance, highlighting text within a code editor gives users the option to either add the code to the right-hand chat interface or edit directly with a prompt. Likewise, inside of the chat interface, buttons appear for users to reference a file or change the scope of the user’s question within a prompt.
4.3.2. Instruction gap
To help users understand how LLMs interpret language, Cursor allows users to ask models questions about their codebase. Topics for these questions can range from the inner workings of a particular function to the overall flow of an entire project. Since this feature can give answers about model-generated and user-written code, developers can get a better idea of how LLMs interpret their programs“under the hood.” A similar feature that can help decipher how models understand code is prompting the model to generate inline comments. Finally, despite no direct functionality for regenerating output outside of feeding the model the same prompt, there is an option to “rerun without context” to see the effects of adding code or references to other materials alongside a prompt. The lack of regeneration history, however, makes it more complicated for users to map changes in their inputs to changes in model outputs.
@@ 4.3.3. Intentionality Gap
The primary way Cursor aligns user intent with model output is through references of code snippets, files, and documentation. When authoring a prompt, users can attach an existing file of code they’ve written, a function within the code editor, or even documentation of third-party libraries to steer the output of the model. This functionality has many use cases, including prompting the model to adopt the same style or asking the LLM to use a particular method or function from a library. While this feature reduces the amount of effort involved in establishing context with the LLM, this affordance does not inherently give users a mental model of how the LLM works, which means that users may still run into issues. Furthermore, to help evaluate output, user can ask Cursor questions about their codebase. The questions can be about any piece of code within the code editor – either model- or user-generated – to help users understand how to make their prompts more optimal.
@@ 5. Recommendations for Designing Interfaces for Envisioning
In addition to the three interfaces described in detail in the previous section, we conducted a qualitative analysis of 12 systems to identify design patterns that would potentially support the process of envisioning. We only selected tools that were either directly accessible or at least had a video demonstration and accompanying technical description of the features, such as in research papers. For each tool, we identified affordances along the three main operators in generative tasks, namely planning, execution, and evaluation. We then clustered the identified features based on their functional similarity to develop a set of interface design patterns and corresponding tenets for teams building human-LLM systems. We do not claim a comprehensive categorization but provide a starting point for the design of interface affordances for envisioning LLM interactions.
@@
Design Pattern 1 – Visually Track Prompts and Outputs: LLM interactions are often iterative and take a trial-and-error approach. We observed that some tools provided users with a visual interface for capturing their prompting and divergent thinking through alternative pathways. A popular choice is node-based interfaces, which allow users to visualize multiple different outputs (nodes) and trace how they are connected (edges). As mentioned before, in Spellburst (Angert et al., 2023), each node represents a sketch, and edges showcase different iterations of the sketch through merging, diffing, and other semantic operators. In PromptChainer (Wu et al., 2022a), a system that helps users chain together LLM prompts for complex tasks, each node represents an individual prompt and corresponding output while each edge represents the use of the output as context for another prompt. Such external representations of their thought process not only lower cognitive load but also help users better engage in more deliberate prompt authoring. For instance, users can readily see if their previous line of thinking was fruitful or make adjustments to subsequent prompts based on prior outputs. These affordances correspond to the tenet that: Users need guidance navigating the cognitive task space for prompt authoring, discerning which paths lead to desired or poor outcomes [T1].
@@
Design Pattern 2 – Suggest Ideas for Prompting: Many systems proactively offer users prompt suggestions. Some of them are aimed at assisting users who may not be as familiar with LLM-enabled interfaces, while others serve as ideation partners in the cognitive task workflow. In addition to providing ideas for prompts, examples and suggestions emphasize the importance of clarity and precision in language. If these suggestions come from examples where the model is fine-tuned, they can also help better align a user’s intentions with model behavior. For example, ChatGPT (OpenAI, 2023a) provides standalone prompt examples to showcase its utility. On the other hand, Notion (Notion, 2023), which is a tool for knowledge management, not only offers possible ways to use AI to improve your writing (i.e., “Change tone”): it also gives suggestions on how to prompt an LLM to steer these improvements in a certain direction (i.e., “Professional,” “Casual,” “Straightforward”). These patterns lead to two Tenets, namely: LLMs can serve as cognitive partners in task formulation [T2] , and A focus on clear, precise written language will help to bridge the gap between human intention and LLM output [T3].
@@
Design Pattern 3 - Provide Multiple Outputs: Rather than generate one output based on a user prompt, LLM systems may provide numerous outputs. This can be achieved by setting the temperature of the model – a parameter that dictates randomness – greater than $0$ and giving the model the same prompt or explicitly asking the model to give more than one example. This feature allows users to view multiple options to see which best fits their intentions. Furthermore, providing multiple outputs helps users link the effects of changes in prompts to changes in the final output of the model. Some systems also support grouping and clustering model outputs to make this process easier. For instance, BotDesigner (Zamfirescu-Pereira et al., 2023b) lets users manually assign a tag to model outputs, while Sensecape (Suh et al., 2023) groups relevant topics together semantically based on a high-level topic. These features support the tenet that LLMs should support users through their divergent thinking strategies [T4].
@@
Design Pattern 4 - Make the Output Explainable: Some systems prompt LLMs to explain their outputs or make them more interpretable. This design pattern allows users to better understand how LLMs interpret certain prompts and makes model outputs easier to use for manual editing. How this technique is applied in practice can differ depending on the task domain. Replit (Replit, 2023), a browser-based code editor, has an AI assistant named Ghostwriter that generates in-line comments within its code responses. Another code editor, Cursor (Anysphere, 2023), does not always provide code comments but does allow users to ask LLMs about the code they generate. In contrast, Sensecape (Suh et al., 2023), which is designed for exploration and sensemaking, prompts an LLM to return a response at different levels of detail, such as through summaries and keywords. These features help users address their intentionality gaps and better assess the model output. This pattern supports the tenet An error in human-LLM interaction is not just a user error or LLM failure but signals a breakdown in the distributed cognitive system that requires collaborative repair [T5]. However, in designing for explainability and drawing causal inferences between prompt inputs and outputs, design should account for users’ overreliance on explanations without careful validation (Fok and Weld, 2023).
@@
Design Pattern 5 - Use domain-specific prompting strategies: Outside of standard prompt engineering techniques, most systems use a custom prompting strategy depending on their task. These methods help steer the outputs from LLMs into something usable for the end goal while also minimizing the output ambiguity that may arise in trying different prompts. As an example, Graphologue (Jiang et al., 2023), which is designed to turn text-based responses from LLMs into graphical diagrams, uses prompting techniques to have models annotate entities and relationships within their outputs to create diagrams in real-time. Coding Steps (Kazemitabaar et al., 2023), a web-based application to help novices learn Python, prompts models with static examples, then user code, then the user prompt, to ensure that the level of output is appropriate for beginners. These strategies allow designers to implement conceptual tasks for users and consequently allow them to build task-specific system mental models. The corresponding tenet is that, Users favor working with a system mental model leading to actions when working within a defined task domain [T6].
@@
Design Pattern 6 - Allow manual control of output: Many systems afford users the opportunity to manually edit the outputs and interactions with LLMs. Since many LLM-enabled systems are built for exploration and ideation, direct manipulation can help users better incorporate their values and intentions into the model. Oftentimes, manual editing is introduced when one output serves as input to another LLM. For instance, while LIDA (Dibia, 2023), a tool for generating visualizations and infographics, prompts an LLM to output goals for dataset exploration, users are also allowed to enter their own goals and adjust the model’s suggestions. Likewise, Mirror (Xu et al., 2023) – an NLI for data querying and summarization – gives users the ability to edit the SQL queries generated by a pre-trained code model to add human expertise. These features align with the tenet, If tasks are well-defined, people prefer dedicated interfaces over dynamic interfaces [T7].